OpenStudy (anonymous):

An equation is shown below: 5(2x - 8) + 15 = -15 Write the steps you will use to solve the equation and explain each step.

OpenStudy (anonymous):

i got thiss

OpenStudy (anonymous):

haha thanks xD

OpenStudy (anonymous):

distributive property again (mulitply the outside number by the inside ) 5(2x-8)+15=-15 10x-40+15=-15 then combine like terms (-40+15) 10x-25=-15 combine like terms again over the = (-25+25 then -15+25) 25 cancels out 10x=10 then divide by 10 x= 1

OpenStudy (anonymous):

wow thanks!

OpenStudy (anonymous):

i think thats right lol medal me maybe?

OpenStudy (anonymous):

thanks c:

OpenStudy (anonymous):